# UKAN: Unbound Kolmogorov-Arnold Network Accompanied with Accelerated Library

### v2

@@ Abstract
In this work, we present a GPU-accelerated library for the underlying components of Kolmogorov-Arnold Networks (KANs), along with an algorithm to eliminate bounded grids in KANs. The GPU-accelerated library reduces the computational complexity of Basis Spline (B-spline) evaluation by a factor of $\mathcal{O}$(grid size) compared to existing codes, enabling batch computation for large-scale learning. To overcome the limitations of traditional KANs, we introduce Unbounded KANs (UKANs), which eliminate the need for a bounded grid and a fixed number of B-spline coefficients. To do so, we replace the KAN parameters (B-spline coefficients) with a coefficient generator (CG) model. The inputs to the CG model are designed based on the idea of an infinite symmetric grid extending from negative infinity to positive infinity. The positional encoding of grid group, a sequential collection of B-spline grid indexes, is fed into the CG model, and coefficients are consumed by the efficient implementation (matrix representations) of B-spline functions to generate outputs. We perform several experiments on regression, classification, and generative tasks, which are promising. In particular, UKAN does not require data normalization or a bounded domain for evaluation. Additionally, our benchmarking results indicate the superior memory and computational efficiency of our library compared to existing codes.
@@ 1 Background
Neural networks (MLPs) are the workhorse of the current AI and Deep Learning revolution, driving advances in computer vision, language models, computational science, and more recently biology and molecular science LeCun et al. (2015); Goh et al. (2017); Schütt et al. (2018); Pandey et al. (2022); Raissi et al. (2017). The universal approximation theorem guarantees that MLPs with enough parameters can fit any function. The widespread adoption of MLPs across various disciplines has led to the emergence of exciting applications such as ChatGPT in LLM and AlphaFold in protein structure prediction Vaswani et al. (2017); Jumper et al. (2021). However, MLPs suffer from a few drawbacks, particularly generalization for regression tasks.
Recently, the Kolmogorov-Arnold network (KAN) Liu et al. (2024) has gained attention as a promising alternative to traditional MLPs, especially in scientific applications, with various variants currently under development Bozorgasl & Chen (2024); Genet & Inzirillo (2024); Abueidda et al. (2024); Kiamari et al. (2024). The KAN architecture is partially based on the Kolmogorov-Arnold representation theorem Kolmogorov (1961); Braun & Griebel (2009), which states that any multivariate function on a bounded domain can be obtained by a finite composition of continuous univariate functions and summation. Mathematically, this can be represented as:
@@
where $\phi_{p,q}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. For a more detailed explanation of the KAN architecture and mathematics, we refer eager readers to the original KAN paper Liu et al. (2024).
Despite their potential, practical use of KANs is currently hindered by compute and memory inefficiencies, which stems from implementation challenges and the partial reliance on grid update tricks. Without grid update tricks, KAN learning stagnates as inputs into the layer can fall outside of the grid domain fixed at the beginning of training. Additionally, the numerical instability of the grid update implementation prevents the original KAN from fully leveraging the capabilities of graphical processing units (GPUs), which are driving current AI advancements. Another drawback of KAN is its inability to leverage batch computation due to the B-spline evaluation being implemented with a complexity of $\mathcal{O}(kd_{g}d_{in}d_{out})$, where $k$, $d_{g}$, $d_{in}$, and $d_{out}$ are the degree of B-spline and grid, input, and output dimensionalities, respectively. However, this computational complexity issue is resolved in this work for the original KAN. We note that B-splines have been used in other works before KAN to learn activation functions, and our library can also be used to accelerate them Bohra et al. (2020). Furthermore, our library facilitates a rigorous and unbiased comparison between KANs and MLPs in terms of floating-point operations (FLOPs), thereby bridging the gap between the typically naive implementations of KANs and the extensively optimized implementations of MLPs that have been refined over decades Yu et al. (2024).
@@ 2 Algorithm
The solution to compute and memory issues of KANs lies in exploring the local nature of B-splines instead of the recursive formula over the entire grid. In other words, the evaluation of the B-spline function is local and depends on the $k+1$ coefficients, leveraging this observation, we represent the B-spline function with basis matrices Qin (1998) as shown in Equation 2.
$P(u)=U\mathbf{M}P$ (2)
where $u=x/\delta g-\lfloor x/\delta g\rfloor$ and $U$ is the vector of $(1,u,u^{2},...,u^{k})$. $\mathbf{M}$ is basis matrix and $P$ is the vector of B-spline coefficients $(p_{0},p_{1},...,p_{k})$. The basis matrices are obtained by applying recursive B-spline equations, and only depends on the degree of the B-spline function, as commonly used in other disciplines Moradzadeh et al. (2018); Mashayak et al. (2015). For a B-spline of order 3, the B-spline evaluation can be written as,
$P(u)=\begin{bmatrix}1&u&u^{2}&u^{3}\end{bmatrix}\frac{1}{6}\begin{bmatrix}1&4&1&0\\ -3&0&3&0\\ 3&-6&3&0\\ -1&3&-3&1\end{bmatrix}\begin{bmatrix}p_{0}\\ p_{1}\\ p_{2}\\ p_{3}\end{bmatrix}$ (3)
We provide efficient implementations of the above formula using NVIDIA Warp Macklin (2022) in a new library called warpKAN with evaluation complexity of $\mathcal{O}(kd_{in}d_{out})$ along with PyTorch bindings Paszke et al. (2019). This implementation offers both memory and computation efficiency, as described in Table 1. However, the reduction in compute and memory cost of B-spline components does not solve the issue of bounded range of the grid in the original KAN.
Figure 1: The UKAN - model architecture including grid group positional encoding, coefficient-generator MLP, and B-spline function.
@@
To achieve an unbounded domain for the KAN grid, we use an MLP to generate B-spline coefficients, called a coefficient-generator (CG) MLP. As shown in Equation 2, every B-spline of order $k$, needs $k+1$ coefficients ($K=k+1$). One approach is to predict coefficient by calling $K$ times same MLP for $K$ adjacent grid indexes, however, in our experimentation we notice this algorithm fails.
To address this issue, we group $K$ adjacent grid indexes together, called a group index, and use the CG model to predict $K$ coefficients for each group index. The selection process works as follows: we first concatenate the $K$ coefficients predicted by the previous group index and the one grid index belongs to, resulting in a sequence of $2K$ coefficients. We then select the $K$ coefficients starting from index $i=g_{id}\mod K$ to $i+K$ ($g_{id}$ is the grid index). This ensures that we obtain the correct $K$ coefficients for each grid index, while minimizing the number of MLP and embedding calls. In other words, instead of embedding every grid index we embed every grid group index i.e. the collection of $K$ adjacent grid indices. The input to the CG MLP is obtained by concatenating the embedding of the feature index and the sinusoidal positional encoding of the group index as used in transformers Vaswani et al. (2017). The architecture of UKAN is depicted in Figure 1.
3 Experiment
We perform several experiments to benchmark the performance of current implementations and capabilities of UKAN in regression, classification, and other tasks.
3.1 Benchmarking
In Figures 2 and 3, we benchmark warp- and torch-KAN in different setups, i.e.: varying grid size and order of the B-spline to evaluate the compute and memory improvements achieved in the current work and the accompanied library. Particularly, in Figure 2, we compare compute cost of the forward & backward passes, and their sum for a single layer KAN [32, 32] with different B-spline orders. During this comparison grid size and batch size of 64 and $2^{16}$ are used and all results for warpKAN are normalized with respect to the PyTorch implementation of the original paper. The results indicate a performance improvement of 5.5-15x which increases as order of B-spline increases.
Figure 2: Performance benchmarking of current and naive implementation of KAN [32, 32] with different B-spline orders. We compared compute time of warp- versus torch-KAN for a single layer with different B-spline orders, (32-dimensional input and output and grid size of 64 on NVIDIA A6000 GPU and batch size of $2^{16}$). We find that warpKAN is on average 10x faster and up to 15x faster for higher order B-splines.
In Figure 3, we benchmark the KAN [32, 32] with B-spline order of 3 on different grid sizes. A batch size of $2^{17}$ is used for this experiment, we observed that torchKAN (naive implementation of B-spline) runs out of memory for a grid size equal and larger than 256, while warpKAN can reach grid sizes of up to $2^{18}$, more than 1000x larger. By reducing the computational and memory costs of B-spline evaluation, KANs can become a more viable option for large-scale applications, paving the way for their widespread adoption beyond toy datasets.
Figure 3: Performance benchmarking of current and naive implementation of KAN [32, 32] with different grid sizes. We compared compute time of warp KAN versus torch KAN with order 3 B-spline and 32-dimensional input and output over different number of grids on NVIDIA A6000 GPU and batch size of $2^{17}$. Torch KAN rans out of memory due to implementation limitations beyond 128, while warp KAN can reach up to $2^{18}$ grids, translating to 1000x larger grid size. We find that warpKAN is on average 12x faster and up to 24x faster for larger grid sizes.

### v4

# Want to train KANS at scale? Now UKAN!

@@ Abstract
Kolmogorov–Arnold Networks (KANs) have recently emerged as a powerful alternative to traditional multilayer perceptrons. However, their reliance on predefined, bounded grids restricts their ability to approximate functions on unbounded domains. To address this, we present Unbounded Kolmogorov–Arnold Networks (UKANs), a method that removes the need for bounded grids in traditional Kolmogorov–Arnold Networks (KANs). The key innovation of this method is a coefficient-generator (CG) model that produces, on the fly, only the B-spline coefficients required locally on an unbounded symmetric grid. UKANs couple multilayer perceptrons with KANs by feeding the positional encoding of grid groups into the CG model, enabling function approximation on unbounded domains without requiring data normalization. To reduce the computational cost of both UKANs and KANs, we introduce a GPU-accelerated library that lowers B-spline evaluation complexity by a factor proportional to the grid size, enabling large-scale learning by leveraging efficient memory management, in line with recent software advances such as FlashAttention and FlashFFTConv. Performance benchmarking confirms the superior memory and computational efficiency of our accelerated KAN (warpKAN), and UKANs, showing a $3-30\times$ speed-up and up to $1000\times$ memory reduction compared to vanilla KANs. Experiments on regression, classification, and generative tasks demonstrate the effectiveness of UKANs to match or surpass KAN accuracy. Finally, we use both accelerated KAN and UKAN in a molecular property prediction task, establishing the feasibility of large-scale end-to-end training with our optimized implementation.
@@ 1 Background
Neural networks (MLPs) are the workhorse of the current AI and Deep Learning revolution, driving advances in computer vision, language models, computational science, and more recently biology and molecular science LeCun et al. (2015); Goh et al. (2017); Schütt et al. (2018); Pandey et al. (2022); Raissi et al. (2017). The universal approximation theorem guarantees that MLPs with enough parameters can fit any function. The widespread adoption of MLPs across various disciplines has led to the emergence of exciting applications such as ChatGPT in the large language model (LLM) domain, and AlphaFold in protein structure prediction Vaswani et al. (2017); Jumper et al. (2021). However, MLPs suffer from a few drawbacks, particularly generalization for regression tasks.
Recently, the Kolmogorov-Arnold networks (KANs) Liu et al. (2024) have gained attention as a promising alternative to traditional MLPs, especially in scientific applications, with novel variants currently under development Bozorgasl & Chen (2024); Genet & Inzirillo (2024); Abueidda et al. (2024); Kiamari et al. (2024); Somvanshi et al. (2024). The KAN architecture is partially based on the Kolmogorov-Arnold representation theorem Kolmogorov (1961); Braun & Griebel (2009), which states that any multivariate function on a bounded domain can be obtained by a finite composition of continuous univariate functions and summation. Mathematically, this can be represented as:
@@
where $\phi_{p,q}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. Eq. 1 implies that we can approximate any function by summation of univariate functions. A B-spline curve is one of the best methods to parameterize any univariate function by learning the coefficients of the B-spline basis function.
To extend the KAN architecture beyond what is described by Eq. 1, Liu et al. (2024) proposes a new way to build the KAN computational graph with $L$ layers. They assume Eq. 1 is a 2-layer KAN with $n$, $2n$, and 1 nodes. The generalization for $L$ layers starts by defining $n_{l}$ nodes $\forall l=0,1,...,L$, where $n_{l}$ is the number of nodes in the $l^{th}$ layer of the computational graph.
The activation function between node $i$ in layer $l$, ($l$,ii) and node $j$ in layer $l+1$, ($l+1$, $j$) is denoted by $\phi_{l,j,i}$; the activation value of node $(l+1,j)$ is obtained by summing all incoming post-activation values $\phi_{l,j,i}(x_{l,i})$,
$x_{l+1,j}=\sum_{i=1}^{n_{l}}\phi_{l,j,i}(x_{l,i})$ (2)
In total, there are $n_{l}n_{l+1}$ activation values and connections between layer $l$ and layer $l+1$. For an input $\mathbf{x}\in\mathbb{R}^{n_{0}}$, a general KAN network can be written as a composition of $L$ layers,
$\operatorname{KAN}(\mathbf{x})=\bigl(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{1}\circ\Phi_{0}\bigr)\,\mathbf{x}.$ (3)
where $\Phi_{l}$ is the matrix function of shape $n_{l+1}\times n_{l}$ with element $(j,i)$ corresponding to $\phi_{l,j,i}$ activation function. In practice, $\phi(x)$ is the sum of the basis function $b(x)$ (MLP with activation function) and the spline function,
$\phi(x)=w_{b}b(x)+w_{s}\operatorname{spline}(x)$ (4)
where $\operatorname{spline}(x)$ is parameterized as a linear combination of B-splines such that
$\operatorname{spline}(x)=\sum_{i}c_{i}B_{i}(x)$ (5)
where $c_{i}$ are learned parameters. A B-spline of order $k+1$ is a collection of piecewise polynomial functions $B_{i,k+1}(t)$ of degree $k$. The locations where these piecewise polynomials connect to each other are known as knots. Given $m+1$ knot values with a uniqueness constraint on $B_{i,k+1}$, we have,
$B_{i,k+1}(t)=\begin{cases}\text{non-zero},&\text{if }t_{i}\leq t<t_{i+k+1}\\[4.0pt] 0,&\text{otherwise }\end{cases}$ (6)
The Cox–de Boor formula recursively builds B-spline of order $k$ using,
$B_{i,k}(t)=\frac{t-t_{i}}{\,t_{i+k}-t_{i}\,}\,B_{i,k-1}(t)+\frac{t_{i+k+1}-t}{\,t_{i+k+1}-t_{i+1}\,}\,B_{i+1,k-1}(t),$ (7)
where $B_{i,0}(t)=\mathbf{1}_{[t_{i},\,t_{i+1})}(t)$.
KANs are promising, but practical use is hampered by compute- and memory-inefficient implementations and reliance on periodic grid updates. In naïve GPU code, evaluating B-spline bases via the Cox–de Boor recursion redundantly recomputes along the entire knot vector (effectively repeating work $k+1$ times per evaluation) and inflates both memory traffic and FLOPs. Many implementations also materialize all intermediate basis states, which depresses batch sizes and GPU utilization. Algorithmically, most public code evaluates B-spline bases in time $\mathcal{O}\!\left(k\,d_{g}\,d_{\mathrm{in}}\,d_{\mathrm{out}}\right)$, where $k$ is the B-spline degree, $d_{g}$ the number of knots (we use “grid” for the sorted knot vector), and $d_{\mathrm{in}}(=n_{l})$, $d_{\mathrm{out}}(=n_{l+1})$ the input and output dimensions. Training can also stall when inputs drift outside the initialized knot support, where the B-spline basis vanishes; while Liu et al. (2024) advocates periodic grid updates, this procedure is numerically brittle and ill-suited to reliable, high-throughput GPU execution.
This work removes these bottlenecks in the original KAN by enabling true batched evaluation and eliminating unnecessary intermediates, substantially reducing both memory footprint and FLOPs. Because B-splines also underlie pre-KAN learnable activations, the same kernels accelerate those methods as well.Bohra et al. (2020) Our implementation facilitates a fair comparison between MLPs and KANs; correcting prior apples-to-oranges comparisons that pitted highly optimized MLPs against vanilla implementation of KANsYu et al. (2024). As recent ML systems work has shown (e.g., FlashAttention/FlashFFTConv Dao et al. (2022); Dao (2023); Shah et al. (2024); Fu et al. (2023)), careful software design unlocks scale; we follow the same philosophy for KAN-style architectures. Finally, we introduce UKAN to eliminate grid updates during training and present experiments comparing MLP, KAN, and UKAN across a range of settings.
@@ 2 Algorithm
Our remedy for KAN’s compute and memory overhead is to exploit the compact support of B-splines rather than evaluating the Cox-de Boor recurrence over the entire knot vector. A degree-k basis $B_{i,k}$ is nonzero only on $t_{i},t_{i+k+1})$. Leveraging this observation, we represent the B-spline function with basis matrices Qin (1998) as shown in Eq. 8.
$\operatorname{spline}(u)=U\mathbf{M}C$ (8)
where $u$ is $\frac{x-t_{i}}{t_{i+1}-t_{i}}$, where $i=\lfloor{\frac{x-t_{0}}{h}}\rfloor$ and $h$ is the distance between two adjacent grid points ($h:=t_{i+1}-t_{i}$). $U$ is the vector of $(1,u,u^{2},...,u^{k})$. $\mathbf{M}$ is basis matrix and $C$ is the vector of B-spline coefficients $(c_{0},c_{1},...,c_{k})$. The basis matrices are obtained by applying recursive B-spline equations, and only depends on the degree of the B-spline function (see Appendix A.1 for cubic B-spline basis matrix representation).
We provide efficient implementations of the above formula using NVIDIA Warp Macklin (2022) in a new library called warpKAN with evaluation complexity of $\mathcal{O}(kd_{in}d_{out})$ along with PyTorch bindings Paszke et al. (2019). This implementation offers both memory and computational efficiency, as described in Table 1. However, the reduction in compute and memory cost of B-spline components does not solve the issue of the bounded range of the grid in the original KAN.
@@
We achieve an unbounded grid by generating B-spline coefficients with a coefficient-generator (CG) MLP, akin to Hyena’s filter generation Poli et al. (2023). For a degree-$k$ spline, each evaluation needs $K=k+1$ coefficients (Eq. 8). A naïve approach—calling the same MLP $K$ times for the $K$ adjacent grid indices performed poorly in our experiments, likely due to missing joint conditioning across those $K$ coefficients. Instead, we partition the uniform grid $\{t_{j}=j\,h\ :\forall j\in\{...,-2,-1,0,1,2,$...\}\} into groups of $K$ consecutive cells. For any $x$, define the cell index $i(x)=\lfloor x/h\rfloor$, the group index
$g=\Big\lfloor\tfrac{i(x)}{K}\Big\rfloor=\Big\lfloor\tfrac{x}{K\,h}\Big\rfloor,$ (9)
and define the within-group offset $r=i(x)\bmod K$. The CG-MLP takes as input the concatenation of (i) an embedding of the feature index and (ii) a sinusoidal positional encoding of the group index $g$ (as in Transformers Vaswani et al. (2017)), and outputs a vector $C_{g}\in\mathbb{R}^{K}$ of coefficients for group $g$. To ensure the correct sliding window across group boundaries, we concatenate the previous and current outputs $C_{g-1};\,C_{g}]\in\mathbb{R}^{2K}$ and select the $K$ coefficients
$\big[C_{g-1};\,C_{g}\big]_{r:\,r+K-1},$ (10)
which are then used in the basis-matrix evaluation of Eq. 8. This grouping yields an unbounded, index-conditioned parameterization with fewer MLP calls, while preserving the exact $K$-wide local stencil required by B-splines. The general architecture of UKAN is shown in Figure 1. After training, one can materialize all coefficients on any finite interval and store them as KAN parameters, retaining interpretability and enabling the usual symbolic-regression or pruning workflows, which is one of the strong features of KAN architecture (see Appendix A.2)
Figure 1: The UKAN model architecture including grid group positional encoding, coefficient-generator MLP, and B-spline function.
3 Experiments
In this section, we benchmark the performance of warpKAN for various B-spline orders and grid sizes. Subsequently, we showcase the capabilities of UKAN across a wide-range of tasks commonly encountered in machine learning: classification, regression, approximation, generation, and a real-world application to drug discovery.
3.1 Performance Benchmarking
In Figure 2, we benchmark warpKAN against torchKAN under two settings: (a) varying the B-spline order and (b) varying the number of knots (“grid size”). We report the speedup for the forward and backward passes, as well as their sum, for a single layer $\mathrm{KAN}(32\!\to\!32)$. For panel (a), we use a grid size of 64 and a batch size of $2^{16}$; for panel (b), we fix the B-spline order to 3 and use a batch size of $2^{17}$. All warpKAN results are normalized to the PyTorch implementation from the original paper. Across B-spline orders, warpKAN is $5.5\!\times$$15\!\times$ faster, with speedup increasing with the order. Across grid sizes, warpKAN averages about $12\!\times$ and reaches up to $24\!\times$ speedup for larger grids. Moreover, torchKAN runs out of memory for grids $\geq 256$, while warpKAN scales to grid sizes up to $2^{18}$; over $1000\!\times$ larger. Notation: $\mathrm{Layer}(a\!\to\!b)$ denotes input and output dimensions.
Figure 2: warpKAN vs. torchKAN. (a) Increasing B-spline order yields larger speedups ($5.5\!\times$$15\!\times$). (b) Increasing grid size shows average $\sim\!12\!\times$ and up to $24\!\times$ speedup; torchKAN hits OOM $\geq 256$ knots while warpKAN scales to $2^{18}$. All values are normalized to the public PyTorch implementation.
From Figure 2, the naïve PyTorch implementation clearly under-utilizes the GPU. Further gains are expected by (i) exploiting Tensor Cores, (ii) fusing kernels to cut launch overhead and off-chip traffic, and (iii) other common techniques in CUDA programming. These optimizations should shift the workload toward the roofline limits and substantially increase throughput, mirroring the trajectory of the FlashAttention family Dao et al. (2022); Dao (2023); Shah et al. (2024).
We quantify how close our implementation is to the hardware limits using a roofline model. The Speed-of-Light (SoL) time, $t_{\mathrm{SoL}}$, corresponds to the limiting factor between compute (FLOP/s) and memory bandwidth (B/s), and thus represents the best achievable runtime on the hardware. Based on this definition, we construct compute and memory models of the B-spline below.
Table 2: B-spline forward SoL and runtime vs. grid size on A6000 (FP32). $t_{meas}$ is our empirical runtime, the remaining columns are theoretical upper and lower bounds. Setup: $N{=}2^{17}$, $d_{\text{in}}{=}d_{\text{out}}{=}32$, order k=3k{=}3 (K=4K{=}4), $P_{\mathrm{FLOP}}{=}38.7$ TFLOP/s, $P_{\mathrm{BW}}{=}768$ GB/s.
$G$ $t_{\text{meas}}$ (ms) $t_{\text{SoL}}^{\text{best}}$ (ms) $t_{\text{SoL}}^{\text{worst}}$ (ms) $t_{\text{compute}}$ (ms) $t_{\text{mem}}^{\text{best}}$ (ms) $t_{\text{mem}}^{\text{worst}}$ (ms)
032 4.9489 0.1387 2.8399 0.1387 0.0439 2.8399
064 5.1764 0.1387 2.8399 0.1387 0.0441 2.8399
128 5.3669 0.1387 2.8399 0.1387 0.0444 2.8399
256 5.5305 0.1387 2.8399 0.1387 0.0451 2.8399
512 5.7400 0.1387 2.8399 0.1387 0.0464 2.8399
1024 5.9188 0.1387 2.8399 0.1387 0.0492 2.8399
Compute model
Let $N$ be the batch size, $d_{\mathrm{in}}$ and $d_{\mathrm{out}}$ the input/output dims, $k$ the B-spline order, K=k+1K=k{+}1, and $G$ the number of grid positions (knots) per $(i,o)$ pair. [^1] Our kernels use a local matrix form; $G$ denotes the global grid size before local selection. We write the per-(input,output) multiply–add count as a $k$-dependent constant $c_{k}$; in our implementation we use $c_{k}\approx 2K(K{+}1)$ (the factor 2 accounts for fused multiply–adds counted as two FLOPs). Then the total work is
$F_{\mathrm{fwd}}\;\approx\;N\cdot d_{\mathrm{in}}\cdot d_{\mathrm{out}}\cdot c_{k}.$ (11)
Memory models
We separate optimistic (best-case) and pessimistic (worst-case) coefficient reuse (no-cache and cached). We assume contiguous reads of inputs and writes of outputs; coefficients are indexed locally ($K$ per sample) but stored globally ((G+K)(G{+}K) per $(i,o)$, including padding).
Best-case reuse (stream all unique coefficients once across the batch):
$B_{\mathrm{best}}\;\approx\;s\Big[N\,(d_{\mathrm{in}}+d_{\mathrm{out}})\;+\;d_{\mathrm{in}}d_{\mathrm{out}}\,(G{+}K)\Big],\qquad s=4~\text{bytes (FP32)}.$ (12)
Worst-case reuse (fetch local $K$ coefficients per sample):
$B_{\mathrm{worst}}\;\approx\;s\Big[N\,(d_{\mathrm{in}}+d_{\mathrm{out}})\;+\;N\,d_{\mathrm{in}}d_{\mathrm{out}}\,K\Big].$ (13)
The SoL times are then
$t_{\mathrm{SoL}}^{\mathrm{best}}=\max\!\left(\frac{F_{\mathrm{fwd}}}{P_{\mathrm{FLOP}}},\frac{B_{\mathrm{best}}}{P_{\mathrm{BW}}}\right),\qquad t_{\mathrm{SoL}}^{\mathrm{worst}}=\max\!\left(\frac{F_{\mathrm{fwd}}}{P_{\mathrm{FLOP}}},\frac{B_{\mathrm{worst}}}{P_{\mathrm{BW}}}\right).$ (14)
We perform experiments for different configurations on the A6000 GPU in FP32, and report results in Table2. We time forward calls with CUDA synchronization before and after, discarding warm-up. We use fixed shapes $(N,d_{\mathrm{in}},d_{\mathrm{out}},k)=(2^{17},32,32,3)$, report the average per-call time, and compute $F_{\mathrm{fwd}}$ via equation 11 with $c_{k}{=}2K(K{+}1)$. We report both $t_{\mathrm{SoL}}^{\mathrm{best}}$ and $t_{\mathrm{SoL}}^{\mathrm{worst}}$ using equation 12–equation 13. As shown in Table 2, there are still opportunities to improve runtime of B-spline used in KANs and many other workloads. The future work will try to close the gap between runtime and SoL time. [^2] We do not provide a full roofline breakdown for UKAN, as its cost is dominated by the CG network. Our ongoing work explores a fused implementation that co-schedules the CG MLP with spline evaluation and accumulation in a single kernel; so that coefficients are produced and consumed without intermediate HBM round-trips. These experiments are in progress. The runtime of vanilla KAN implementations can be estimated by multiplying the speedup values reported in Figure 2 with the measured $t_{\mathrm{meas}}$ from Table 2, provided the configuration does not run out of memory.